OpenStudy (anonymous):

Gene bought 27 comics and 12 magazines for $126. If 3 comics cost as much as 2 magazines find the cost of each comics. Thank you,

OpenStudy (lgbasallote):

let x = comics and y = magazine 27x + 12y = 126 3x = 2y using systems of equations... 27+ 12y = 126 (3x - 2y = 0) 6 so.. 27 + 12y = 126 18x - 12y = 0 45x = 126 x = 2.8 there are 3 comics

OpenStudy (anonymous):

we know there are 3 comics but how much is each comic?

OpenStudy (lgbasallote):

i'm sorry. My conclusion was wrong. $3 is each comic. 3 is not the number of comics; 27 is

OpenStudy (anonymous):

they do not equal $126 if each comic is $3, I think this is incorrect. Any other way to solve?